If 2*3=9, 5*3=18, 6*4=30, then 7*4=? …......................

Grade:6

2 Answers

Arun
4 years ago
2*3 + 3 = 9
5*3 + 3 = 18
6*4 + 6 = 30
7*4 + 7 = 35
 
Hope it helps
In case of any difficulty, please feel free to ask
Thanks and regards
Vikas TU
4 years ago
Dear student 
The above ans is incorrect 
2×3=6+3= 9
5×3=15+3=18
6×4=24+3+3= 30
7×4+3+3=34
Hope it helps